x=0,1,2,3,4
y=0.5,1.5,2.5,3.5,4.5
write a rule for the function?

Answer by stanbon(75887)

x=0,1,2,3,4
y=0.5,1.5,2.5,3.5,4.5
write a rule for the function?
--------------------
There is a consistent difference between the x-values and the y-values
so the function is linear.
---
slope = (1.5 - 0.5)/(1-0) = 1
intercept is the y-value that goes with x=0; int = 1/2
---------------------
Y = x+(1/2)
=====================
Cheers,
Stan H>